If the population consists of numbers 1, 2 ,3, 4, 5 and 6, find the mean of the samples of size of 3. Construct a sampling distribution and make a histogram for the sample distribution.


1
Expert's answer
2021-05-07T10:24:49-0400

Mean of The samples is shown in the table below.


Mean of sample mean "=\dfrac{\sum x}{n}=\dfrac{30.63}{10}=3.063"


Sampling distribution of given data is-




Histogram for the sample distribution is-